OpenStudy (anonymous):

If A= 30 and B =90 what is P(A|B) P(B|A)

OpenStudy (anonymous):

Use the following formulas: \[\huge P(A|B) = \frac{P(A \cap B)}{P(B)}\] \[\huge P(B|A) = \frac{P(A \cap B)}{P(A)}\] And in case of Independent Events: \[\huge P(A \cap B) = P(A).P(B)\]